# HOIDiNi: Human-Object Interaction through Diffusion Noise Optimization

### v1

@@ Abstract.
We present HOIDiNi, a text-driven diffusion framework for synthesizing realistic and plausible human-object interaction (HOI). HOI generation is extremely challenging since it induces strict contact accuracies alongside a diverse motion manifold. While current literature trades off between realism and physical correctness, HOIDiNi optimizes directly in the noise space of a pretrained diffusion model using Diffusion Noise Optimization (DNO), achieving both. This is made feasible thanks to our observation that the problem can be separated into two phases: an object-centric phase, primarily making discrete choices of hand-object contact locations, and a human-centric phase that refines the full-body motion to realize this blueprint. This structured approach allows for precise hand-object contact without compromising motion naturalness. Quantitative, qualitative, and subjective evaluations on the GRAB dataset alone clearly indicate HOIDiNi outperforms prior works and baselines in contact accuracy, physical validity, and overall quality. Our results demonstrate the ability to generate complex, controllable interactions, including grasping, placing, and full-body coordination, driven solely by textual prompts. https://hoidini.github.io.
Human motion, Animation, Motion synthesis, Computer Graphics.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
Figure 1. HOIDiNi generates human-object interactions from text descriptions and object geometry, integrated here into a 3D scene from Jay-Artist (2012).
[…]
fig:teaser
1. Introduction
@@
Human-object interaction (HOI) lies at the core of many everyday tasks, such as frying an egg or drinking a cup of water, with crucial applications to any digital human-like agent. Even though it is ubiquitous, this central capability still eludes modern motion generation modeling techniques. This is because HOI modeling requires millimeters-level accuracy to avoid noticeable artifacts, but tackles a diverse motion space, rich with nuanced human behavior. Indeed, when applying traditional generation techniques to the rather limited HOI data available, results often exhibit physical artifacts like inter-object penetration, floating, or implausible grasps — even when restricted to hand-only scenarios  (Huang et al., 2025; Zhang et al., 2025a; Li et al., 2024a).
@@
To facilitate desired accuracy while maintaining plausibility, most recent literature employs generation guidance (Peng et al., 2023; Diller and Dai, 2024; Zhang et al., 2025b) or post-generation optimization (Wu et al., 2024; Ghosh et al., 2023). While this can improve physical correctness, both of these approaches adhere to fine-grained contact requirements by pulling the motion off the human manifold, on the account of realism.
@@
In this work, we present Human-Object Interaction through Diffusion Noise optimization (HOIDiNi), a text-driven diffusion framework that satisfies the tight constraints of HOI while remaining on the manifold of realistic human motion. We address this challenge using an optimization strategy that, by design, preserves the learned motion distribution. Diffusion Noise Optimization (DNO) (Karunratanakul et al., 2024), a test-time sampling method that traverses the noise space of a pretrained diffusion model to steer generation toward desired losses. Originally applied to control free-form motion synthesis, DNO proves to be a natural fit for HOI when carefully adapted to the structure and demands of the task.
@@
A central challenge in the first phase is determining which locations on the hands should make contact with the object and where. Typically in prior works (Zhang et al., 2025b, 2024a), this is done using nearest-neighbor heuristics, but this approach is brittle, especially for small or thin objects, and highly sensitive to initialization. Instead, we explicitly predict contact pairs between the hand and object surfaces, and optimize them jointly with the object’s 6-DoF trajectory. The DNO objective enforces semantically meaningful placement while preventing interpenetration with supporting surfaces.
@@
The second phase then optimizes the full-body motion, including the fingers, conditioned on the object trajectory and contact pairs produced in the first phase. In this case, DNO helps satisfy these contacts without penetrating the object. Throughout, the DNO process keeps the samples close to the motion manifold, ensuring realism.
@@
Quantitative evaluations on the GRAB dataset (Taheri et al., 2020a) demonstrate that HOIDiNi outperforms prior baselines in both interaction accuracy and motion realism, as measured by contact precision, physical validity, and proximity to the human motion manifold. Subjectively, a user study indicates dramatic preference to our resulting motions compared to competing literature and baselines. Qualitatively, we showcase a range of full-scene interactions, including object grasping and placement, all driven by textual instructions, and trained on the GRAB dataset alone. These results highlight HOIDiNi’s ability to synthesize complex, controllable, and physically plausible HOI behaviors.
@@
Figure 2. System Overview. HOIDiNi generates Human-object Interaction (HOI) motions according to a text prompt, a mesh describing the object, and the dimensions of the surface on which it is placed. The Object-Centric Phase generates the object motion and its contact points with the hands, then the Human-Centric Phase follows and generates the full human motion: body and fingers, adhering to the constraints implied by the previous phase. Both phases use CPHOI, a pre-trained diffusion model that learned the human-object joint distribution. They apply Diffusion Noise Optimization (DNO) (Karunratanakul et al., 2024) to fulfill the two sets of loss functions ($\mathcal{L}_{\text{Object}}$ and $\mathcal{L}_{\text{Human}}$) without deviating from the learned distribution.
2. Related Work
@@
Controlled Motion Synthesis. Current motion synthesis methods increasingly focus on controllability. TEMOS (Petrovich et al., 2022) and MotionCLIP (Tevet et al., 2022) addressed text-to-motion synthesis using a Transformer VAE (Vaswani et al., 2017; Kingma and Welling, 2013). T2M (Guo et al., 2022a), T2M-GPT (Zhang et al., 2023), and MoMask (Guo et al., 2024) adopt a VQ-VAE (Van Den Oord et al., 2017) to quantize motion and generate it sequentially in the latent space, conditioned on text. MDM (Tevet et al., 2023) and MoFusion (Dabral et al., 2023) introduced the denoising diffusion framework (Ho et al., 2020a) to motion synthesis, demonstrating its effectiveness across multimodal tasks such as action-, text-, and music-to-motion (Tseng et al., 2023).
@@
The diffusion paradigm enables diverse control mechanisms: PriorMDM (Shafir et al., 2024), CondMDI (Cohan et al., 2024), GMD (Karunratanakul et al., 2023), and OmniControl (Xie et al., 2023) combine temporal conditioning with Classifier-guidance (Dhariwal and Nichol, 2021) to achieve joint-level control. MoMo (Raab et al., 2024) demonstrated motion transfer through attention injection, while LoRA-MDM (Sawdayee et al., 2025) employed Low-Rank Adaptation (Hu et al., 2022) for motion stylization.
CAMDM (Chen et al., 2024) and A-MDM (Shi et al., 2024) accelerated sampling by introducing autoregressive motion diffusion. CLoSD (Tevet et al., 2025) further integrated autoregressive diffusion into a physics-based simulation framework for object interaction.
@@
Human Object Interaction. SAGA (Wu et al., 2022) adopt a two-stage approach: they first generate a static target frame and then synthesize the in-between motion using a VAE-based decoder. GOAL (Taheri et al., 2022) uses a similar framework with additional optimization refrainment to algin the motion and object. IMoS(Ghosh et al., 2023) employs dual-stream networks to autoregressively generate arm and body motions, followed by an optimization step to align the object for plausible hand-object interaction. TOHO(Li et al., 2024b) first predicts the object’s final position, then generates human poses to grasp the object and fills in the motion trajectory using an implicit neural representation.
@@
Recent works have increasingly adopted diffusion models for HOI generation. OOD-HOI (Zhang et al., 2024b) employs a dual-branch reciprocal diffusion model to synthesize motions, followed by optimization similar to IMoS for refinement. Both HOI-Diff (Peng et al., 2023) and DiffGrasp (Zhang et al., 2025b) uses classifier guidance for fine-grained corrections of grasping motion. Whereas HOI-Diff utilize affordance-guided classifier-based diffusion, DiffGrasp used goal functions to correct the motion. OMOMO(Li et al., 2023) predicts the joint pathologies of the hand with a diffusion model conditioned on the motion of the object, which then guides a second diffusion model for the full-body synthesis, but does not handle the motion of the fingers. Similarly, BimArt(Zhang et al., 2024a) conditions contact generation on object trajectories, using the predicted contact sequence to guide body motion. Both methods reduce task complexity by generating only partial motion (hands or body), enabling more tractable modeling at the cost of full-scene coherence.
Recently, CLoSD (Tevet et al., 2025) and Wu et al. (2024) suggested applying physical trackers on top of the generated motion, which increases grasp accuracy at the cost of motion realism.
3. Preliminaries
Denoising diffusion models.
Denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020b) define a forward Markov process $\{x_{t}\}_{t=0}^{T}$ that progressively adds Gaussian noise to a clean data sample $x_{0}\sim p_{\text{data}}(x_{0})$. At each timestep $t$, the noising process is defined as: $q(x_{t}\mid x_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}x_{t-1},(1-\alpha_{t})I),$ where $\alpha_{t}\in(0,1)$ is a predefined noise schedule. As $t$ increases, the data distribution converges to an isotropic Gaussian, i.e., $x_{T}\sim\mathcal{N}(0,I)$.
In the reverse process, the model learns to gradually denoise a sample from this prior distribution back to the data manifold. In the conditional setting, the goal is to model $p(x_{0}|c)$, where $c$ denotes some conditioning signal (e.g., text or pose). While DDPMs originally parameterized the model to predict the noise term $\epsilon_{t}$, we adopt the alternative parameterization proposed by MDM (Tevet et al., 2023) which yields better results for motion synthesis. Here, the model directly predicts the clean sample $\hat{x}_{0}$ at each step. This leads to the simplified training objective:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{x_{0}\sim p(x_{0}\mid c),\,t\sim[1,T]}\left[|x_{0}-\hat{x}_{0}|_{2}^{2}\right].$
Diffusion noise optimization (DNO)
@@ 3. Preliminaries
A common approach to constrain a data sample $x\!\sim\!X$ is to directly optimize it via $x^{*}=\arg\min_{x}\mathcal{L}(x)$, where $\mathcal{L}$ encodes task-specific objectives. In the context of HOI, such post-hoc optimization has been widely used (Ghosh et al., 2023; Zhang et al., 2024a; Paschalidis et al., 2024), but it lacks guarantees that $x^{*}$ remains within the data distribution $X$, often resulting in unrealistic outputs. A more robust strategy is to optimize in a latent space $z\sim Z$, assuming $x=D(z)$ for some decoder $D$. The optimization becomes:
@@
where $\mathcal{R}(z)$ encourages $z\sim Z$, thereby keeping the final output $x^{*}=D(z^{*})$ on-manifold. This technique is commonly employed with VAEs (Holden et al., 2016; Pavlakos et al., 2019a) and GANs (Karras et al., 2020a; Patashnik et al., 2021) across both image and 3D domains.
@@
Where the final output will be $x^{*}=\text{ODE}(G,x_{T}^{*})$. Here, $\mathcal{R}_{\text{decorr}}$ is a decorrelation regularizer (Karras et al., 2020b) that encourages $x_{T}$ to remain within the Gaussian prior. HOIDiNi extends this formulation with a two-phase DNO strategy tailored for HOI, enabling precise contact control while preserving motion realism.
Figure 3. CPHOI Diffusion Model. CPHOI autoregressively predicts the next motion segment $s^{n}$ from the previous one $s^{n-1}$. The figure illustrates a single diffusion step, where the model denoises $s_{t}^{n}$ to predict $\hat{s}_{0}^{n}$. It jointly generates human and object motions, along with dynamic contact points, conditioned on the object’s geometry and a text description of the interaction.
4. Method
@@
At inference, we employ a two-phase Diffusion Noise Optimization (DNO) (Karunratanakul et al., 2024) strategy tailored to HOI (4.3). The first, object-centric, phase optimizes the object’s trajectory and contact pairs based on scene constraints such as placement and support. The second, human-centric, phase completes the full-body motion, including hand articulation, to fulfill the previously determined contact goals while maintaining realistic posture and avoiding collisions. This structure allows us to satisfy complex physical constraints while remaining within the learned motion distribution.
Figure 4. Additional Results of human-object interactions generated by our method across diverse prompts. For instance, “taking a picture with a camera” yields a semantically appropriate two-handed pose. Motions are both physically plausible and aligned with the prompts.
4.1. Data Representation
@@
Contact Pairs Sequence
The Contact Pairs sequence is central to our method, generated during the first, object-centric, phase and serving as the conditioning signal in the second, human-centric phase. It is defined as:
$\left\{\left[\left(i^{\text{Body}}_{f,a},\ i^{\text{Object}}_{f,a}\right)\right]_{a=1}^{A_{f}}\right\}_{f=1}^{F}$
where each pair $(i^{\text{Body}}_{f,a},\ i^{\text{Object}}_{f,a})$ denotes a contact between a vertex on the hand and a vertex on the object at frame $f$, and $A_{f}$ denotes the number of contact pairs at that frame.
Directly generating index pairs via diffusion is challenging. To address this, we adopt a more learnable representation. We define a fixed anchor set $\mathcal{A}$, consisting of a subset of MANO (Romero et al., 2017a) hand vertices located on the palm (Figure 9). At each frame $f$, and for each anchor $a\in\mathcal{A}$, a binary variable $b_{a}$ indicates whether the anchor is in contact. A corresponding position $p_{a}\in\mathbb{R}^{3}$ specifies the contact location on the object surface in its rest pose. This yields the contact representation:
resulting in a per-frame contact feature of dimension $(3+1)\times|\mathcal{A}|$.
Human Motion Representation
We adopt a variant of the widely used HumanML3D (Guo et al., 2022b) representation to encode human motion. The per-frame human feature is defined as:
$F_{H}=\left[r_{z}^{H},\dot{r}_{x}^{H},\dot{r}_{y}^{H},\dot{\alpha}^{H},\theta^{H},j^{H}\right]$
where $r_{z}^{H}$ is the vertical root height, $(\dot{r}_{x}^{H},\dot{r}_{y}^{H})$ denotes the planar root velocity, $\dot{\alpha}^{H}$ is the angular root velocity, $\theta^{H}$ contains the SMPL-X pose parameters, and $j^{H}$ denotes the relative 3D joint positions.
The 52 relevant joints from the SMPL-X model are used, including both body and hand joints. Unlike the original representation, we directly employ the SMPL-X pose parameters $\theta^{H}$, allowing us to extract the human mesh in a fully differentiable way. This property is essential for enabling backpropagation during the diffusion noise optimization process.
Object Motion Representation
The per-frame object’s pose is, $F_{O}=\left[\theta^{O},r^{O},\dot{r}^{O}\right]$, where $\theta^{O}$ denotes the object’s global rotation represented in Cont6d format, $r^{O}$ is the object’s global translation, and $\dot{r}^{O}$ is its linear velocity. Together, these parameters define the object’s 6DoF trajectory.
Final representation
Combining the representations of our three data components, we end up with a feature representation for each frame with the form $\mathbf{F}=[F_{CP},F_{H},F_{O}]$. It is illustrated at the bottom of Figure 3
4.2. CPHOI Diffusion Model

### v2

@@ Abstract
We present HOIDiNi, a text-driven diffusion framework for synthesizing realistic and plausible human-object interaction (HOI). HOI generation is extremely challenging since it induces strict contact accuracies alongside a diverse motion manifold. While current literature trades off between realistic motions and accurate contacts, HOIDiNi optimizes directly in the noise space of a pretrained diffusion model using Diffusion Noise Optimization (DNO), achieving both. This is made feasible thanks to our observation that the problem can be separated into two phases: an object-centric phase, primarily making discrete choices of hand-object contact locations, and a human-centric phase that refines the full-body motion to realize this blueprint. This structured approach allows for precise hand-object contact without compromising motion naturalness. Quantitative, qualitative, and subjective evaluations on the GRAB and OMOMO datasets clearly indicate HOIDiNi outperforms prior works and baselines in contact accuracy, visual plausibility, and overall quality. Our results demonstrate the ability to generate complex, controllable interactions, including grasping, placing, and full-body coordination, driven solely by textual prompts. https://hoidini.github.io/
1 Introduction
@@
Human-object interaction (HOI) lies at the core of many everyday tasks, such as frying an egg or drinking a cup of water, with crucial applications to any digital human-like agent. Even though it is ubiquitous, this central capability still eludes modern motion generation modeling techniques. This is because HOI modeling requires millimeters-level accuracy to avoid noticeable artifacts, but tackles a diverse motion space, rich with nuanced human behavior.
Indeed, when applying traditional generation techniques to the rather limited HOI data available, results often exhibit physical artifacts like inter-object penetration, floating, or implausible grasps, even when restricted to hand-only scenarios  (Huang et al., 2025; Zhang et al., 2025a; Li et al., 2024a).
@@
To facilitate desired accuracy while maintaining plausibility, most recent literature employs generation guidance (Peng et al., 2023; Diller & Dai, 2024; Zhang et al., 2025b) or post-generation optimization (Wu et al., 2024; Ghosh et al., 2023). While this can improve physical correctness, both of these approaches adhere to fine-grained contact requirements by pulling the motion off the human manifold, on the account of realism.
@@
In this work, we present Human-Object Interaction through Diffusion Noise optimization (HOIDiNi), a text-driven diffusion framework that satisfies the tight constraints of HOI while remaining on the manifold of realistic human motion. We address this challenge using an optimization strategy that, by design, preserves the learned motion distribution: Diffusion Noise Optimization (DNO) (Karunratanakul et al., 2024), a test-time sampling method that traverses the noise space of a pretrained diffusion model to steer generation toward desired losses. Originally applied to control free-form motion synthesis, DNO proves to be a natural fit for HOI when carefully adapted to the structure and demands of the task.
@@
A central challenge in the first phase is determining which locations on the hands should make contact with the object and where. Typically in prior works Zhang et al. (2025b; 2024a), this is done using nearest-neighbor heuristics, but this approach is brittle, especially for small or thin objects, and highly sensitive to initialization. Instead, we explicitly predict contact pairs between the hand and object surfaces, and optimize them jointly with the object’s 6-DoF trajectory. The DNO objective enforces semantically meaningful placement while preventing interpenetration with supporting surfaces.
@@
After this outline of object motion and contact locations is determined and fixed, the second phase then optimizes the full-body motion, including the fingers, conditioned on the object trajectory and contact pairs. In phase case, DNO helps satisfy these contacts without penetrating the object. Throughout, the DNO process keeps the samples close to the motion manifold, ensuring realism.
@@
Quantitative evaluations on the GRAB (Taheri et al., 2020b) and OMOMO (Li et al., 2023) datasets demonstrate that HOIDiNi outperforms prior baselines in both interaction accuracy and motion realism, as measured by contact precision, physical validity, and proximity to the human motion manifold. Subjectively, a user study indicates dramatic preference to our resulting motions compared to competing literature and baselines. Qualitatively, we showcase a range of full-scene interactions, including object grasping and placement, all driven by textual instructions,trained on a single dataset. These results highlight HOIDiNi’s ability to synthesize complex, controllable, and visually plausible HOI behaviors.
Figure 1: HOIDiNi generates human-object interactions from text descriptions and object geometry, integrated here into a 3D scene from Jay-Artist (2012).
@@
Figure 2: System Overview. HOIDiNi generates Human-object Interaction (HOI) motions according to a text prompt, a mesh describing the object, and the occupied volume in the scene, by optimizing the diffusion noise. The Object-Centric Phase generates the object motion and its contact points with the hands (CPCP and $O$), then the Human-Centric Phase follows and generates the full human motion($H$): body and fingers, adhering to the constraints implied by the previous phase. Both phases use CPHOI, a pre-trained diffusion model that learned the human-object joint distribution. We apply Diffusion Noise Optimization (DNO) (Karunratanakul et al., 2024) to fulfill the two sets of loss functions ($\mathcal{L}_{\text{Object}}$ and $\mathcal{L}_{\text{Human}}$) without deviating from the learned distribution.
2 Related Work
@@
Controlled Motion Synthesis. Current motion synthesis methods increasingly focus on controllability. TEMOS (Petrovich et al., 2022) and MotionCLIP (Tevet et al., 2022) addressed text-to-motion synthesis using a Transformer VAE (Vaswani et al., 2017; Kingma & Welling, 2013). T2M (Guo et al., 2022b), T2M-GPT (Zhang et al., 2023), and MoMask (Guo et al., 2024) adopt a VQ-VAE (Van Den Oord et al., 2017) to quantize motion and generate it sequentially in the latent space, conditioned on text. MDM (Tevet et al., 2023) and MoFusion (Dabral et al., 2023) introduced the denoising diffusion framework (Ho et al., 2020b) to motion synthesis, demonstrating its effectiveness across multimodal tasks such as action-, text-, and music-to-motion (Tseng et al., 2023).
@@
The diffusion paradigm enables diverse control mechanisms: PriorMDM (Shafir et al., 2024), CondMDI (Cohan et al., 2024), GMD (Karunratanakul et al., 2023), and OmniControl (Xie et al., 2023) combine temporal conditioning with Classifier-guidance (Dhariwal & Nichol, 2021) to achieve joint-level control. MoMo (Raab et al., 2024) demonstrated motion transfer through attention injection, while LoRA-MDM (Sawdayee et al., 2025) employed Low-Rank Adaptation (Hu et al., 2022) for motion stylization. CAMDM (Chen et al., 2024) and A-MDM (Shi et al., 2024) accelerated sampling by introducing autoregressive motion diffusion. CLoSD (Tevet et al., 2025) further integrated autoregressive diffusion into a physics-based simulation framework for object interaction.
@@
Human-Object Interaction. Early HOI methods generate motions in stages: SAGA (Wu et al., 2022) first predicts a static target frame, then interpolates motion via a VAE decoder; GOAL (Taheri et al., 2022) adds optimization to align motion and object; IMoS (Ghosh et al., 2023) uses dual-stream autoregressive networks for arm and body motions followed by object alignment; TOHO (Li et al., 2024b) predicts the object’s final position, generates grasping poses, and fills the trajectory with an implicit representation. Diffusion-based methods have recently gained traction: OOD-HOI (Zhang et al., 2024b) uses a dual-branch reciprocal diffusion model with IMoS-style refinement; HOI-Diff (Peng et al., 2023), CHOIS (Li et al., 2024a), and DiffGrasp (Zhang et al., 2025b) apply classifier guidance, with HOI-Diff using affordances and CHOIS and DiffGrasp goal functions; OMOMO (Li et al., 2023) models hand-object paths first, then full-body motion; BimArt (Zhang et al., 2024a) conditions contact generation on object trajectories to guide body motion. Both methods reduce task complexity by generating only partial motion (hands or body), enabling more tractable modeling at the cost of full-scene coherence. Finally, CLoSD (Tevet et al., 2025) and Wu et al. (2024) apply physical trackers atop generated motion, improving grasp accuracy at the cost of realism.
Concurrent to this work, CoDa (Pi et al., 2025) applied Diffusion Noise Optimization with separate hand and body models, thus not learning the joint Hand–Body distribution.
Figure 3: Contact Pairs. CPHOI predicts precise, semantically meaningful contact points between the hand and object. Each contact pair is visualized with matching colored spheres.
3 Preliminaries
Denoising diffusion models. Denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020a) define a forward Markov process $\{x_{t}\}_{t=0}^{T}$ that progressively adds Gaussian noise to a data sample $x_{0}\sim p_{\text{data}}(x_{0})$: $q(x_{t}\mid x_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}x_{t-1},(1-\alpha_{t})I),$ with $\alpha_{t}\in(0,1)$. As $t$ increases, $x_{T}$ approaches $\mathcal{N}(0,I)$. The reverse process learns to denoise back to $x_{0}$, optionally conditioned on $c$ (e.g., text or pose). Unlike the original DDPM that predicts noise $\epsilon_{t}$, we follow MDM (Tevet et al., 2023) and predict $\hat{x}_{0}$ directly, yielding the training objective: $\mathcal{L}_{\text{simple}}=\mathbb{E}_{x_{0}\sim p(x_{0}\mid c),\,t\sim[1,T]}\left[|x_{0}-\hat{x}_{0}|_{2}^{2}\right].$
@@ 3 Preliminaries
Diffusion noise optimization (DNO). A common approach to constrain a data sample $x\sim X$ is to directly optimize it via $x^{*}=\arg\min_{x}\mathcal{L}(x)$, where $\mathcal{L}$ encodes task-specific objectives. In the context of HOI, such post-hoc optimization has been widely used (Ghosh et al., 2023; Zhang et al., 2024a; Paschalidis et al., 2024), but it lacks guarantees that $x^{*}$ remains within the data distribution $X$, often resulting in unrealistic outputs. A more robust strategy is to optimize in a latent space $z\sim Z$, assuming $x=D(z)$ for some decoder $D$. The optimization becomes:
@@
where $\mathcal{R}(z)$ encourages $z\sim Z$, thereby keeping the final output $x^{*}=D(z^{*})$ on-manifold. This technique is commonly employed with VAEs (Holden et al., 2016; Pavlakos et al., 2019b) and GANs (Karras et al., 2020b; Patashnik et al., 2021) across both image and 3D domains.
@@
Where the final output will be $x^{*}=\text{ODE}(G,x_{T}^{*})$. Here, $\mathcal{R}_{\text{decorr}}$ is a decorrelation regularizer (Karras et al., 2020a) that encourages $x_{T}$ to remain within the Gaussian prior. HOIDiNi extends this formulation with a two-phase DNO strategy tailored for HOI, enabling precise contact control while preserving motion realism.
Figure 4: CPHOI Diffusion Model. CPHOI autoregressively predicts the next motion segment $s^{n}$ from the previous one $s^{n-1}$. The figure illustrates a single diffusion step, where the model denoises $s_{t}^{n}$ to predict $\hat{s}_{0}^{n}$. It jointly generates human and object motions, along with dynamic contact points, conditioned on the object’s geometry and a text description of the interaction.
4 Method
@@
At inference, we employ a two-phase Diffusion Noise Optimization (DNO) (Karunratanakul et al., 2024) strategy tailored to HOI (4.3). The first, object-centric, phase optimizes the object's trajectory and contact pairs based on scene constraints such as placement and support. The second, human-centric, phase completes the full-body motion, including hand articulation, to fulfill the previously determined contact goals while maintaining realistic posture and avoiding collisions. This structure allows us to satisfy complex physical constraints while remaining within the learned motion distribution.
Figure 5: Qualitative Results of human-object interactions generated by our method across diverse prompts. For instance, “taking a picture with a camera” yields a semantically appropriate two-handed pose. Motions are both visually plausible and aligned with the prompts.
4.1 Data Representation
@@
Contact Pairs Sequence. The Contact Pairs sequence is central to our method, generated during the first, object-centric, phase and serve as conditioning for the second phase. Unlike previous approaches, that snap the given or predicted object trajectory to the hands using simple heuristics, we adopt a learned approach to semantically predict these contacts. We define a discrete set of Anchor points $\mathcal{A}$ on the fingers and palms, denoting potential contacts between the hands and the object (See Appendix). At each frame $f$, and for each anchor $a\in\mathcal{A}$, we represent and predict contacts using a binary variable $b_{a}$, indicating whether the anchor is in contact, and a corresponding position $p_{a}\in\mathbb{R}^{3}$, specifying the contact location on the object surface in its rest pose. This yields the contact representation: $F_{CP}=\left[p_{1},\dots,p_{|\mathcal{A}|},b_{1},\dots,b_{|\mathcal{A}|}\right]$ of dimension $(3+1)\times|\mathcal{A}|$ per-frame in the sequence.
Human Motion Representation. We adopt a variant of the widely used HumanML3D (Guo et al., 2022a) representation to encode human motion. The per-frame human feature is defined as: $F_{H}=\left[r_{z}^{H},\dot{r}_{x}^{H},\dot{r}_{y}^{H},\dot{\alpha}^{H},\theta^{H},j^{H}\right]$ where $r_{z}^{H}$ is the vertical root height, $(\dot{r}_{x}^{H},\dot{r}_{y}^{H})$ denotes the planar root velocity, $\dot{\alpha}^{H}$ is the angular root velocity, $\theta^{H}$ contains the SMPL-X pose parameters, and $j^{H}$ denotes the relative 3D joint positions.
The 52 relevant joints from the SMPL-X model are used, including both body and hand joint rotations (but not shape). Unlike the original representation, we directly employ the SMPL-X pose parameters $\theta^{H}$, allowing us to extract the human mesh in a fully differentiable manner. This property is essential for enabling backpropagation during the diffusion noise optimization process.
Object Motion Representation. The per-frame object’s pose is, $F_{O}=\left[\theta^{O},r^{O},\dot{r}^{O}\right]$, where $\theta^{O}$ denotes the object’s global rotation represented in Cont6d format, $r^{O}$ is the object’s global translation, and $\dot{r}^{O}$ is its linear velocity. Together, these parameters define the object's 6DoF trajectory.
Final representation. Combining the representations of our three data components, we end up with a feature representation for each frame with the form $\mathbf{F}=[F_{CP},F_{H},F_{O}]$ (Figure 4, bottom)
4.2 CPHOI Diffusion Model